OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 1)2 – 16? Identify the x–intercepts by graphing the quadratic equation

OpenStudy (cwrw238):

the y intercept is value of y when x = 0

OpenStudy (cwrw238):

the x intercepts are values when y = 0

OpenStudy (anonymous):

these are my choices: x–intercepts: (3, 0) and (–5, 0) y–intercept: (0, –16) x–intercepts: (–3, 0) and (5, 0) y–intercept: (0, –16) x–intercepts: (3, 0) and (–5, 0) y–intercept: (0, –15) x–intercepts: (–3, 0) and (5, 0) y–intercept: (0, –15)

OpenStudy (cwrw238):

replace x by 0 in the equation enables you to work out y intercept y = (0-1)^2 - 16

OpenStudy (cwrw238):

can you work that out?

OpenStudy (anonymous):

no im still confused

OpenStudy (cwrw238):

y = 1 - 16 (since (0-1)^2 = (-1)^2 = 1)

OpenStudy (cwrw238):

1 - 16 = ?

OpenStudy (anonymous):

-15

OpenStudy (cwrw238):

right - so we now know that the y intercept iS (0,-15) either 3rd or 4th option is correct

OpenStudy (cwrw238):

to find x intercepts we solve the equation (x - 1)^2 - 16 = 0 ( because y = 0 at the x-intercepts) (x - 1)^2 = 16 (adding 16 to both sides of equation) now take square root of both sides x - 1 = 4 also x - 1 = -4 (both 4 and -4 squared = 16) can you continue?

OpenStudy (anonymous):

How would I continue?

OpenStudy (cwrw238):

first x - 1 = 4 so what is x?

OpenStudy (anonymous):

3?

OpenStudy (cwrw238):

no its 5 add 1 to both sides (to get x only on left side) x - 1 + 1 = 4 + 1 x = 5

OpenStudy (cwrw238):

now we need to solve the second equation x - 1 = -4

OpenStudy (anonymous):

-3

OpenStudy (cwrw238):

right so x intercepts = (5,0) and (-3,0) ( note that y = 0 at x-intercepts)

OpenStudy (anonymous):

ahh I see how you did it

OpenStudy (cwrw238):

right - so you now correct option now

OpenStudy (anonymous):

thx

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

graph is below: http://www.wolframalpha.com/input/?i=y++%3D+%28x+-+1%29%5E2+-+16+

OpenStudy (anonymous):

thx!